# Spatial Context-based Self-Supervised Learning for Handwritten Text Recognition

### v1

@@ 1 Introduction
Handwritten text recognition (HTR) is the research area in the field of computer vision whose objective is to transcribe the textual content of a written manuscript into a digital machine-readable format [73]. This field not only plays a key role in the current digital era of handwriting by electronic means (such as tablets) [11], but is also of paramount relevance for the preservation, indexing and dissemination of historical manuscripts that exist solely in a physical format [56].
HTR has developed considerably over the last decade owing to the emergence of Deep Learning [57], which has greatly increased its performance. However, in order to attain competitive results, these solutions usually require large volumes of manually-labelled data, which is the principal bottleneck of this method. One means by which to alleviate this problem, Self-Supervised Learning (SSL), has recently gained considerable attention from the research community [61]. SSL employs what is termed as a pretext task to leverage collections of unlabelled data for the training of neural models in order to obtain descriptive and intelligible representations [8], thus reducing the need for large amounts of labelled data [4].
The pretext tasks can be framed in different categories according to their working principle [34, 61], with the following being some of the main existing families: (i) image generation strategies [63, 46], which focus on recovering the original distribution of the data from defined distortions or corruptions; (ii) contrastive learning methods [60, 33], whose objective is to learn representative and discernible codifications of the data, and (iii) spatial context methods [27, 58], which focus on either estimating geometric transformations performed on the data [27]i.e. spatial information—or inferring the disposition of a set of patches in which the image is divided as a puzzle [21, 58]i.e. structural information.
Nevertheless, the development of SSL in computer vision has predominantly focused on image classification [45], while other domains such as HTR have garnered noticeably less attention [36]. In this respect, while there are several research efforts that explore ad-hoc adaptations from strategies used in classification tasks [3, 81, 50] or new methods particularly devised for these cases [49, 66], there is a remarkable scarcity of proposals, thus hindering the development of the HTR field.
@@
This work was motivated by the relevance of spatial information in HTR tasks [49], together with the conceptual simplicity and competitive performance of spatial context-based SSL strategies, and focuses on the adaptation of a reference set of such approaches to the HTR field. More precisely, we propose two alternatives that enable the use of these SSL strategies in HTR scenarios (see Fig. 1): (i) adapting the HTR task to the original (or classification-based) SSL method, namely input adaptation, and (ii) adapting the particular SSL method to the characteristics of HTR task, namely method adaptation. These proposals are comprehensively evaluated by considering several reference HTR corpora and are compared to the reference strategies followed in the field. The results obtained show that the proposed SSL strategies not only provide better recognition rates than exclusively relying on annotated data but also outperform the existing state-of-the-art methods for SSL pre-training of HTR.
The remainder of the work is structured as follows: Section 2 contextualizes the current work within the related literature of the field, while Section 3 introduces the strategies devised for spatial context-based SSL for HTR; Section 4 presents the experimental set-up considered, and Section 5 presents and analyses the results obtained. Finally, Section 6 provides the conclusion found after carrying out this work.
@@ 2.1 Handwritten Text Recognition
While HTR initially relied on Hidden Markov Models [40, 23, 68, 71], advances in Machine Learning, and specifically in Deep Learning, favoured the use of Recurrent Neural Networks (RNNs) owing to their capacity to model sequential data [28, 17, 42, 31]. In particular, the use of bidirectional Long Short-Term Memory (LSTMs) networks with Connectionist Temporal Classification (CTC) [28] has been the state of the art in HTR competitions for decades [5, 70, 69, 2]. Despite this prevalence of CTC, attention-based encoder-decoder approaches [7] have recently gained popularity owing to their competitive results [37, 41, 1, 18]. For a detailed review of this field, see the work of Michael et al. [52], which provides a comprehensive study of the various sequence-to-sequence approaches for HTR.
As has occurred in many other areas, there has been a growing interest in the use of more parallelizable architectures such as the Transformer [72] by adapting the work of [22] to the HTR field [10, 64, 67, 26, 20, 65]. HTR has benefited from this adaptation, either in isolation with an encoder-decoder [43, 54, 55, 74] or in combination with the CTC objective function [75, 19, 9]. The work of Diaz et al. [20] explores universal architectures for text recognition, concluding that a CNN backbone with a Transformer encoder, a CTC-based decoder, plus an explicit language model, is the most effective strategy to date. Regardless of this progress, however, the need for large labelled corpora as a pre-training strategy in Transformer-based models has become noticeable,[43, 20, 47, 79, 48]. This has led the SSL field to attract more interest in the HTR field owing to the usual scarcity of annotated data for particular domains.
@@ 2.2 Self-Supervised Text Recognition
SSL is currently a promising alternative for the pre-training of text recognition models, especially when addressing scenarios in which there is a scarcity of labelled data [4]. In this context, the first work that tackles this issue is SeqCLR [3], which adapts the SimCLR contrastive learning method [13] to the sequential nature of text recognition. Please note that, while this work is tested for both HTR and Scene Text Recognition (STR), subsequent SSL proposals have been specifically devised in order to address the nuances of each scenario. We now introduce some of the most relevant approaches in each case.
With regard to STR, PerSEC [44] is a competitive example of a contrastive framework based on extracting low-level stroke and high-level semantic contextual spaces simultaneously via a hierarchical learning approach. SimAN [49], on the contrary, is framed within a generative strategy based on recovering the original format of an image crop by means of another non-overlapping crop of the same datum. This work highlights the close relationship between HTR and spatial context information. Other noticeable generative approaches rely on Mask Image Modelling for the pre-training of full Transformer architectures, as occurs in [50, 66, 77].
In the case of HTR, the current state of the art is defined by ChaCo [81], which takes advantage of Momentum Contrast [29, 15] to pre-train a visual feature extractor by considering a Character Unit Cropping module, which ensures that two crops of the same HTR image have an overlapping part. The same authors also show that the pre-training of only the visual feature extractor and not the sequence one improves the state of the art in semi-supervised scenarios, and this is also supported by the work of Luo et al. [49].
@@ 2.3 Spatial context-based methods
Image data inherently encodes spatial information, which can be exploited in order to pre-train recognition models. According to this, SSL approaches based on this principle constitute rather simple yet effective pretext tasks. However, despite being largely explored in classification frameworks, they have rarely been studied in text recognition. We now present the two principal families of approaches that follow this principle, in addition to highlighting the main works related to them.
In terms of geometric transformations, the work of Gidaris et. al. [27] introduced the concept of predicting rotation alteration on image data as an SSL strategy. Its success in terms of both performance and simplicity led other works to explore this principle in greater depth in order to further exploit this premise [59, 80, 14, 24, 76].
Since the relative position among patches of an image encompasses rich spatial and contextual information, there are several pretext tasks that exploit this premise, namely jigsaw puzzle strategies. Some of the most common approaches rely on predicting the relative positions of two patches of an image [21] or recognising the order of a shuffled sequence of patches of an image [58, 12, 62, 53, 78].
Spatial context is rarely exploited in image-based text recognition tasks. Nevertheless, the competitive results that are obtained by employing these strategies for classification tasks, together with the strong relationship between spatial information and HTR—owing to stroke and font style—, suggest that, if adequately adapted, text recognition models could benefit from pre-training methods of this nature. Note that our work not only seeks to understand and unravel the existing spatial context SSL methods, but also presents innovative methods that specifically leverage and capitalize on this spatial context, thus providing new perspectives and advancements within HTR.
@@ 3 Methodology
This section introduces the general recognition framework considered, after which the proposed spatial context-based SSL methods are presented and described. Fig. 2 depicts a graphical example of the proposed pretext tasks for HTR.
Figure 2: Pre-training pipeline for the different pretext tasks proposed in the work. An image is transformed according to the method and the subsequent layers handle the task. Note that Rotation, Flip and the Jigsaw puzzle predict a single class from the sets presented, while the Sorting method predicts a sequence.
@@ 3.1 Encoder-decoder architecture
In this paper, a state-of-the-art encoder-decoder pipeline is adopted for text recognition, as shown in Fig. 3. The encoder is responsible for extracting the visual and sequential characteristics of an image, while the decoder is responsible for transforming the representation into the output text.
Figure 3: Encoder-decoder pipeline. The encoder is composed of a CNN that extracts the visual features for a given image. After a map-to-sequence operation, an RNN extracts the sequential features in order to provide the final representation. The decoder, which may in this case be either Connectionist Temporal Classification or a Transformer Decoder, transforms the representation into the output text.
Encoder. Given an image $I\in\mathbb{R}^{c\times H\times W}$ as input, where $c$ is the input channels, $H$ is the height and $W$ is the width, the first goes through a Convolutional Neural Network (CNN) $\text{cnn}:\mathbb{R}^{c\times H\times W}\rightarrow\mathbb{R}^{c^{\prime}\times H^{\prime}\times W^{\prime}}$, which is responsible for extracting visual features, where $c^{\prime}$ is the number of feature maps of size $H^{\prime}\times W^{\prime}$. A map-to-sequence operation $m:\mathbb{R}^{c^{\prime}\times H^{\prime}\times W^{\prime}}\rightarrow\mathbb{R}^{W^{\prime}\times\left(c^{\prime}\cdot H^{\prime}\right)}$ then reshapes the features into a sequence of frames. Finally, a Recurrent Neural Network (RNN) $\text{rnn}:\mathbb{R}^{W^{\prime}\times\left(c^{\prime}\cdot H^{\prime}\right)}\rightarrow\mathbb{R}^{W^{\prime}\times T}$ is responsible for extracting sequential features, where $T$ is the output size that leads to the final representation $S=[s_{1},s_{2},...,s_{W^{\prime}}]$. The encoder $f(\cdot)$ can, therefore, be expressed as $f(\cdot)=\text{rnn}(m(\text{cnn}(\cdot)))$
Decoder. Given the representation $S$, in this work two different text decoders are used: (i) a CTC decoder that separately decodes each frame of $S$, and (ii) a Transformer Decoder (TD), which is responsible for autoregressively producing the text characters on the basis of the representation $S$ and the previously generated output.
@@ 3.2 SSL pipeline
This section presents the proposed adaptation of the main existing spatial context-based pretext tasks to HTR scenarios. Note that, for each pretext task, we present two possible means of enabling the use of these techniques in HTR scenarios: (i) adapting the data to a classification-oriented framework (input adaptation), and (ii) modifying the underlying nature of the pretext task in order to allow its used in sequential-based data (method adaptation).
@@ 3.2.1 Geometric transformations
As stated previously, one of the most relevant strategies in terms of geometric transformations is the prediction of rotations when applied to an image [27]. While a transformation of this nature can be applied directly in classification scenarios, the typology of the input in HTR data makes its direct use complicated, since: i) text images are rarely squared, signifying that most rotations alter the dimensionality of the axes, and ii) applying a rotation to such elements drastically changes the semantic meaning of the axes (characteristics and sequential order). We, therefore, now present the two possible adaptations of this technique to HTR scenarios:
  • Rotation (input adaptation). Given that images in HTR rarely depict the same width, a fixed-size crop—for simplicity, we consider that a square crop of dimensions $H\times H$—of the input data must be extracted in order to be able to use this pretext task. This pre-training process can be formally described as follows: given an image $I$, a squared crop is made, obtaining $I_{square}$, after which a rotation is uniformly chosen from the set {0º, 90º, 180º, 270º}, retrieving image $I_{r}$. Finally, the CNN $cnn(\cdot)$ will, together with a linear classification layer $l(\cdot)$, be responsible for solving the pretext task.
@@ 3.2.2 Puzzle transformation
One of the most prominent methods in this family is the so-called jigsaw puzzle, which consists of rearranging a set of jumbled patches from the same image by predicting the index of the correct permutation [58]. Since this proposal was conceived for classification tasks in which images depict the same dimensions, we now present the alternatives conceived in order to enable its use in HTR images with different width dimensions.
  • Jigsaw puzzle (input adaptation). The original method is maintained by extracting a square crop from the HTR image, and this is later processed by employing the jigsaw puzzle strategy. Formally, for a given image $I$, a square crop of $H\times H$ dimensions is obtained from the image, obtaining $I_{square}$; the image is initially reshaped to obtain $I_{square}^{\prime}\in\mathbb{R}^{c\times 150\times 150}$ and then divided into 4 shuffled patches of dimensions $75\times 75$; these patches are further cropped into patches of size $64\times 64$, i.e., $P=[p_{1},p_{2},p_{3},p_{4}]$, where $p_{i}\in\mathbb{R}^{1\times 64\times 64}$; each patch is subsequently processed by the CNN feature extractor, whose results are concatenated and fed into a multi-layer perceptron that estimates the permutation performed.
  • Sorting (method adaptation). The adaptation of the method proposed in this work consists of rearranging a fixed-size non-overlapping number of patches—experimental parameter—into which the HTR image is divided. Formally, given an input image $I$ and a number of patches $M$—in this case, $M\in\{2,3,4\}$—, a list of patches $P=[p_{1},...,p_{M}]$, where $p_{i}\in\mathbb{R}^{1\times H\times\lfloor W/M\rfloor}$, is obtained. $P$ is then randomly unsorted and rejoined to form the image $I_{shuffled}$. Finally, the encoder $f(\cdot)$ predicts the correct order of the patches using an auxiliary decoder.
@@ 4 Experimental setup
This section presents the datasets and figures of merit considered in order to validate the proposals presented, along with the information regarding the neural architectures used and their implementation details.

### v2

@@ 1 Introduction
Handwritten Text Recognition (HTR) is the research area in the field of computer vision whose objective is to transcribe the textual content of a written manuscript into a digital machine-readable format [36]. This area not only plays a key role in the current digital era of handwriting by electronic means (such as tablets) [5], but is also of paramount relevance for the preservation, indexing and dissemination of historical manuscripts that exist solely in a physical format [28].
HTR has developed considerably over the last decade owing to the emergence of Deep Learning [29], which has greatly increased its performance. However, to attain competitive results, these solutions usually require large volumes of manually-labelled data, which constitutes their main bottleneck. One means by which to alleviate this problem, Self-Supervised Learning (SSL), has recently gained considerable attention from the research community [32]. SSL employs a so-called pretext task to leverage collections of unlabelled data for the training of neural models to obtain descriptive and discriminative representations [4], thus reducing the need for large amounts of labelled data.
The pretext tasks can be framed in different categories according to their working principle [18, 32], with the following being some of the main existing families: (i) image generation strategies [24], which focus on recovering the original distribution of the data from defined distortions or corruptions; (ii) discriminative learning methods [17], whose objective is to learn representative and discernible codifications of the data. In the last one, we can find spatial context methods, which focus on either estimating geometric transformations performed on the data [13]i.e. spatial information—or inferring the disposition of a set of patches in which the image is divided as a puzzle [11, 30]i.e. structural information.
Nevertheless, the development of SSL in computer vision has predominantly focused on image classification, whereas other domains such as HTR have garnered noticeably less attention. In this respect, although there are several research efforts that explore ad-hoc adaptations from strategies used in classification tasks [1, 41] or new methods particularly devised for these cases [25, 33], there is a remarkable scarcity of proposals, thus leaving substantial SSL areas unexplored hindering the development of the HTR field.
@@
This work was motivated by the relevance of spatial information in HTR tasks [25] together with the fact that, to our best knowledge, it has not been previously studied, and focuses on the adaptation of a reference set of such approaches to the HTR field. More precisely, we propose two alternatives that enable the use of these SSL strategies in HTR scenarios (see Fig. 1): (i) adapting the HTR task to the original SSL method, namely input adaptation, and (ii) proposing novel spatial context-based SSL methods devised to the characteristics of the HTR task, namely method adaptation. These proposals are comprehensively evaluated by considering several reference HTR corpora and are compared to the reference strategies followed in the field. The results obtained show that the proposed SSL strategies provide state-of-the-art recognition rates compared to reference SSL approaches with less computational requirements, while maintaining conceptual simplicity.
The remainder of the work is structured as follows: Section 2 contextualizes the current work within the related literature of the field, while Section 3 introduces the strategies devised for spatial context-based SSL for HTR; Section 4 presents the experimental set-up considered, and Section 5 presents the results obtained; Section 6 provides an in-depth analysis and discussion of the strengths and limitations of the proposed methods; finally, Section 7 concludes the work.
@@ 2.1 Handwritten Text Recognition
While HTR initially relied on Hidden Markov Models [34], advances in Machine Learning favored the use of Recurrent Neural Networks (RNN) owing to their capacity to model sequential data. In particular, the use of bidirectional Long Short-Term Memory (LSTM) networks with Connectionist Temporal Classification (CTC) [14] has been the state of the art in HTR for years [36]. Despite this prevalence of CTC, attention-based (Att.) autorregresive encoder-decoder approaches [3], have recently gained popularity owing to their competitive results. For a detailed review of HTR approaches, see the work of [27].
As in other areas, there has been a growing interest in the use of more parallelizable architectures such as the Transformer [35] by adapting the work of [12]. HTR has benefited from this adaptation, either in isolation with fully Transformer architectures [22] or in combination with other approaches [10]. Regardless of this progress, the need for large labelled corpora as a pre-training strategy in Transformer-based models has become noticeable [22]. This has led the SSL field to attract more interest in the HTR field owing to the usual scarcity of annotated data.
@@ 2.2 Self-Supervised Text Recognition
SSL represents a promising alternative for pre-training text recognition models, especially when addressing scenarios depicting scarcity of labelled data [2]. In this context, the first work that tackled this issue was SeqCLR [1], which adapted the SimCLR contrastive learning method [7] to the sequential nature of text recognition. Note that, while this work is tested for both HTR and Scene Text Recognition (STR), subsequent SSL proposals have been specifically devised to address the nuances of each scenario. We now introduce some of the most relevant approaches in each case.
With regard to STR, PerSEC [23] is a competitive example of a contrastive framework based on extracting low-level stroke and high-level semantic contextual spaces simultaneously via a hierarchical learning approach. SimAN [25], on the contrary, is framed within a generative strategy based on recovering the original format of an image crop by means of another non-overlapping crop of the same datum. This work highlights the close relationship between HTR and spatial context information. Other noticeable generative approaches rely on Mask Image Modelling (MIM) for the pre-training of full Transformer architectures, as occurs in the works by [33] and [38].
In the context of HTR, the ChaCo [41] and CMT-Co [40] methods constitute representative examples of the contrastive learning paradigm. Both strategies take advantage of Momentum Contrast [16] to pre-train a visual feature extractor, demonstrating that focusing solely on that part rather than the sequential can improve state-of-the-art performances in semi-supervised scenarios, which is also supported by the work of [25]. Regarding MIM, we find Text-DIAE [33], which is based on denoising autoencoders for Transformer architectures.
@@ 2.3 Spatial context-based methods
Image data inherently encodes spatial information, which can be exploited to pre-train recognition models. We now present the two main families of SSL approaches that follow this principle highlighting the most important works related to them.
In terms of geometric transformations, the work of [13] introduced the concept of predicting rotation alteration on image data as an SSL strategy. Its success in terms of both performance and simplicity led other works to further explore this principle in greater depth [31, 8, 37].
Since the relative position among patches of an image encompasses rich spatial and contextual information, there are several pretext tasks that exploit this premise, namely puzzle strategies. Some of the most common approaches rely on predicting the relative positions of two patches of an image [11] or recognising the order of a shuffled sequence of patches of an image [30, 6, 39].
Despite being largely explored in classification frameworks, spatial context-based SSL is rarely exploited in image-based text recognition tasks. Nevertheless, the competitive results that are obtained by employing these strategies for classification tasks, together with the strong relationship between spatial information and HTR—owing to stroke and font style—suggest that, if adequately adapted, text recognition models should benefit from pre-training methods of this nature. Note that, our work not only seeks to understand and unravel the existing spatial context SSL methods, but also presents innovative methods that specifically leverage and capitalize on this spatial context, thus providing new perspectives and advancements within HTR.
@@ 3 Methodology
In this section, we introduce the general recognition framework considered and detail the spatial context-based SSL methods proposed in this work.
@@ 3.1 Encoder-decoder architecture
A state-of-the-art encoder-decoder pipeline is adopted for the text recognition task, as shown in Fig. 2. The encoder is responsible for extracting the visual and sequential characteristics of an image, while the decoder is responsible for transforming the representation into the output text.
Figure 2: Encoder-decoder pipeline. The encoder is composed of a CNN that extracts the visual features for a given image. After a map-to-sequence operation, an RNN extracts the sequential features. The decoder, which may in this case be either CTC or a Transformer Decoder, transforms the representation into the output text.
Encoder. Given an image $I\in\mathbb{R}^{c\times H\times W}$ as input, where $c$ is the input channels, $H$ is the height and $W$ is the width, the first goes through a Convolutional Neural Network (CNN) $\text{cnn}:\mathbb{R}^{c\times H\times W}\rightarrow\mathbb{R}^{c^{\prime}\times H^{\prime}\times W^{\prime}}$, which is responsible for extracting visual features, where $c^{\prime}$ is the number of feature maps of size $H^{\prime}\times W^{\prime}$. A map-to-sequence operation $m:\mathbb{R}^{c^{\prime}\times H^{\prime}\times W^{\prime}}\rightarrow\mathbb{R}^{W^{\prime}\times\left(c^{\prime}\cdot H^{\prime}\right)}$ then reshapes the features into a sequence of frames. Finally, a RNN $\text{rnn}:\mathbb{R}^{W^{\prime}\times\left(c^{\prime}\cdot H^{\prime}\right)}\rightarrow\mathbb{R}^{W^{\prime}\times T}$ is responsible for extracting sequential features, where $T$ is the output size that leads to the final representation $S=[s_{1},s_{2},...,s_{W^{\prime}}]$. The encoder $f(\cdot)$ can, therefore, be expressed as $f(\cdot)=\text{rnn}(m(\text{cnn}(\cdot)))$
Decoder. Given the representation $S$, in this work two different text decoders are used: (i) a CTC decoder that separately decodes each frame of $S$, and (ii) a Transformer Decoder (TD), which is responsible for autoregressively producing the text characters from the representation S and the previously generated output.
@@ 3.2 SSL pipeline
This section presents the proposed adaptations for the main existing spatial context-based pretext tasks to HTR scenarios (see Fig. 3). Note that for each spatial SSL family—i.e. geometric and puzzle—we have chosen the most representative pretext task. Then, for each selected method, we present two possible means of enabling the use of these techniques in HTR scenarios: (i) adapting the data to a classification-oriented framework (input adaptation), and (ii) modifying the underlying nature of the pretext task to allow its used in HTR-based data (method adaptation).
Figure 3: Pre-training pipeline for the different pretext tasks proposed in the work. An image is transformed according to the method and the subsequent layers handle the task. Note that “Rotation”, “Flip” and “Jigsaw” predict a single class from the sets presented, while “Sorting” predicts a sequence.
@@ 3.2.1 Geometric transformations
As stated previously, one of the most relevant strategies in terms of geometric transformations is the prediction of rotations when applied to an image [13]. While a transformation of this nature can be applied directly in classification scenarios, the typology of the input in HTR data makes its direct use complicated, since: i) text images are rarely squared, signifying that most rotations alter the dimensionality of the axes; and ii) applying a rotation to such elements drastically changes the semantic meaning of the axes (characteristics and sequential order). Therefore, we explore two possible adaptations of this technique to HTR scenarios:
  • Rotation (input adaptation). Given that images in HTR rarely depict the same width, a fixed-size crop—for simplicity, we consider that a square crop of dimensions $H\times H$—of the input data must be extracted to enable the use of this pretext task. This pre-training process can be formally described as follows: given an image $I$, a squared crop is made, obtaining $I_{square}$, after which a rotation is uniformly chosen from the set {0º, 90º, 180º, 270º}, retrieving image $I_{r}$. Finally, the CNN $cnn(\cdot)$ will, together with a linear classification layer $l(\cdot)$, be responsible for solving the pretext task.
@@ 3.2.2 Puzzle transformation
One of the most prominent methods in this family is the so-called jigsaw puzzle, which consists of rearranging a set of jumbled patches from the same image by predicting the index of the correct permutation [30]. Since this proposal was conceived for classification tasks in which images depict the same dimensions, we now present the proposals to enable its use in HTR images with different width dimensions.
  • Jigsaw puzzle (input adaptation). The original method is maintained by extracting a square crop from the HTR image, and this is later processed by employing the jigsaw puzzle strategy. Formally, for a given image $I$, a square crop of $H\times H$ dimensions is extracted from the image, obtaining $I_{square}$; the image is initially reshaped to obtain $I_{square}^{\prime}\in\mathbb{R}^{c\times 150\times 150}$ and then divided into 4 shuffled patches of dimensions $75\times 75$; these patches are further cropped into elements of size $64\times 64$, i.e., $P=[p_{1},p_{2},p_{3},p_{4}]$, where $p_{i}\in\mathbb{R}^{1\times 64\times 64}$; each patch is subsequently processed by the CNN feature extractor, whose results are concatenated and fed into a multi-layer perceptron that estimates the permutation performed among the $24$ ($4!$) possibilities.
  • Sorting (method adaptation). The proposed method adaptation consists of rearranging a fixed-size non-overlapping number of patches—experimental parameter—into which the HTR image is divided. Formally, given an input image $I$ and a number of patches $M$—in this case, $M\in\{2,3,4\}$—, a list of patches $P=[p_{1},...,p_{M}]$, where $p_{i}\in\mathbb{R}^{1\times H\times\lfloor W/M\rfloor}$, is obtained. $P$ is then randomly unsorted and rejoined to form the image $I_{shuffled}$. Finally, the encoder $f(\cdot)$ predicts the correct index order of the patches using an auxiliary decoder.
@@ 4 Experimental setup
This section presents the datasets and the metric considered in order to validate the proposals presented, along with the information regarding the neural architectures used and their implementation details.